Gunther invested $10,000 in two mutual funds. One of the funds rose 6 percent in one year and the other rose 9 percent in one year. If Gunthers investment rose a total of $684 in one year, how much did he invest in each mutual fund?
Yes, this is my son's 7th graded algebra homework. He's learning to solve systems of equations so my first inclination is to .06x +.09y=684 and x + y = 10,000

Found 2 solutions by rvquartz, mananth:
Answer by rvquartz(19)

that is correct!

Answer by mananth(16949)

Yes, this is my son's 7th graded algebra homework. He's learning to solve systems of equations so my first inclination is to .06x +.09y=684 and x + y = 10,000
0.06x +0.09y=684 ...........1
x + y = 10,000 ............2
...
multiply (1) by 100
we get
6x+9y=68400................3
Multiply (2) by -6
we get
-6x-6y=-60000.............4
Add (3) & (4)
6x-6x+9y-6y=68,400-60,000
3y=8,400
/3
y=2800
So x= 7,200